# Multitask methods for predicting molecular properties from heterogeneous data

### v1

@@ Appendix B SOAP parameters
Appendix C Global Features
@@
Fig. 8 shows the correlation in kernel outputs corresponding to each pair of molecules for different global featurization strategies. Consistently, in our tests, we find strong correlation (Pearson’s coefficient >0.99)>0.99)> 0.99 ) between the REMatch features and features computed by averaging all local representations. These two strategies are also both positively correlated with the “Average by Species” approach which produces an averaged feature for each element in the system and inserts isolated atoms to represent elements not included in the system, but these correlations are not as strong as that between the REMatch and totally averaged features. We may see this result because the REMatch and totally averaged features share the same dimension, $n$, whereas features that are constructed by species specific averaging have dimension of $2n$ to $4n$ in the cases we considered. Because averaging features is much more computationally efficient than constructing REMatch features, we use this approach in our numerical experiments.

### v2

@@ Appendix C SOAP parameters
Appendix D Global Features
@@
Fig. 9 shows the correlation in kernel outputs corresponding to each pair of molecules for different global featurization strategies. Consistently, in our tests, we find strong correlation (Pearson’s coefficient >0.99)>0.99)> 0.99 ) between the REMatch features and features computed by averaging all local representations. These two strategies are also both positively correlated with the “Average by Species” approach which produces an averaged feature for each element in the system and inserts isolated atoms to represent elements not included in the system, but these correlations are not as strong as that between the REMatch and totally averaged features. We may see this result because the REMatch and totally averaged features share the same dimension, $n$, whereas features that are constructed by species-specific averaging have dimension of $2n$ to $4n$ in the cases we considered. Because averaging features is much more computationally efficient than constructing REMatch features, we use this approach in our numerical experiments.